If the hypotenuse of a right triangle is 2 m more than the longer leg, and the shorter leg is 14 m long, find the length of the longer leg.
Answer by rwm(914)

14^2+l^2=(l+2)^2
l=48 and hypotenuse =50
14^2+48^2=50^2